Please help in all 3 sections!
When a is with a particular wire vibrating frequency of 4 Hz, a transverse wave of wave- length 29 cm is produced. Determine the speed of wave pulses along the wire. Answer in units of m/s.
media%2F498%2F498ed4c6-1fce-4980-b077-8f
media%2Fba7%2Fba75b918-de58-48f1-a540-42
Answer #1

Question 1 :

using v = f\lambda , here v is speed of wave , \lambda is wavelength and f is frequency

so v = 29*10-2*4 = 1.16 m/sec

Question 2 :

wavelength of detected wave = \lambda = v/f

\lambda = 419/(50.7*103) m

\lambda = 8.2643*10-3 m = 0.826 cm

so it can detect minimum 0.826 cm size insect.

Question 3 :

He will detect the flash immediatly.

Distance = speed of sound*time taken by sound

d = 343*4.86 = 1666.98 m